OpenStudy (anonymous):

Fortuna worked 49.5 hours this week at $8.18 an hour. What was her gross pay? Remember that over time worked over 40 hours is paid at time and a half.

OpenStudy (anonymous):

$366.055

OpenStudy (anonymous):

other these $356.98 $359.42 $361.86 $364.30

OpenStudy (anonymous):

Do you mean that those are the multiple choice options.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

I'm sorry, I just realized my mistake. Just give me a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's not coming up right for me. I think this is how you do it: 8.18*40=327.20 8.18+4.09=12.27 12.27*9.5=116.565 116.565+327.20=443.765 I must be wrong though. Try tagging somebody who might know a little more about math.

OpenStudy (anonymous):

@wio, can you try?

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

What was the answer?